OpenStudy (anonymous):

solvign using complete the square: 2x^2-4x-8=0

OpenStudy (anonymous):

Take the 2 out first 2(x^2-2x-4)=0 2(x^2-2x+1-1-4)=0 2((x-1)^2-5)=0 2(x-1)^2-10=0 2(x-1)^2=10 (x-1)^2=5 Can you handle the rest?

OpenStudy (anonymous):

yes thanks!!!!

OpenStudy (anonymous):

wait where did you get the +1-1 in the second step?

OpenStudy (anonymous):

coefficient of x divided by 2 then squared (-2/2)^2=1 you add and subtract that after x

OpenStudy (anonymous):

oh okay thanks!